OpenStudy (anonymous):

Solve the equation 2/5 x − 3 = 13 for x. Please show your work and solution in the space provided.

OpenStudy (campbell_st):

add 3 to both sides, whar do you get..?

OpenStudy (amorfide):

 you want to multiply both sides by 5 to get rid of the fraction this gives 2x-15=65 re arrange to make it 2x=C 2x=80 x=80/2 x=40